Integrate the following with respect to x:
-33(5x4/1+e-x)dx =

Grade:10

2 Answers

Aditya Gupta
4 years ago
hello meenu, let J= -33(5x4/1+e-x)dx …..(1)
now, we know by properties of definite integral that we can replace the fn f(x) inside by f(a+b – x) or f( – 3+3 – x)= f( – x)
so, J= -33(5( – x)4/1+ex)dx
or J= -33(5x4/1+ex)dx ….....(2)
now, write (1) as J= -33(5ex.x4/1+ex)dx …......(3)
add (2) and (3)
2J= -33(5x4/1+ex)*(1+ex)dx = -335x4dx = x^5| – 3 to 3= 2*3^5
or 2J= 486
so that J = 243
kindly approve :))
Vikas TU
4 years ago
Dear student 
For more information please follow this link 
Good Luck